Hula Enterprises is considering a new project to produce solar water heaters. The finance manager wishes to find an appropriate risk adjusted discount rate for the project. The (equity) beta of Hot Water, a firm currently producing solar water heaters, is 1.2. Hot Water has a debt to total value ratio of 0.6. The expected return on the market is 0.11, and the riskfree rate is 0.04. Suppose the corporate tax rate is 38 percent. Assume that debt is riskless throughout this problem. (Round your answers to 2 decimal places. (e.g., 0.16))

a. The expected return on the unlevered equity (return on asset, R0) for the solar water heater project is  %.
b. If Hula is an equity financed firm, the weighted average cost of capital for the project is  %.
c. If Hula has a debt to equity ratio of 1, the weighted average cost of capital for the project is  %.
d. The finance manager believes that the solar water heater project can support 30 cents of debt for every dollar of asset value, i.e., the debt capacity is 30 cents for every dollar of asset value. Hence she is not sure that the debt to equity ratio of 1 used in the weighted average cost of capital calculation is valid. Based on her belief, the appropriate debt ratio to use is  %. The weighted average cost of capital that you will arrive at with this capital structure is  %.
Answer #1

a

D/A = 0.6

D/E = D/(A-D) = 0.6/(1-0.6)=1.5

Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
1.2 = Unlevered Beta*(1+((1-0.38)*(1.5)))
Unlevered Beta = 0.62
As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 4 + 0.62 * (11 - 4)
Expected return% = 8.34 = unlevered cost of equity

b

For all equity firm WACC = unlevered cost of equity = 8.34%

c

Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 8.34+1*(8.34-4)*(1-0.38)
Levered cost of equity = 11.03
D/A = D/(E+D)
D/A = 1/(1+1)
=0.5
Weight of equity = 1-D/A
Weight of equity = 1-0.5
W(E)=0.5
Weight of debt = D/A
Weight of debt = 0.5
W(D)=0.5
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 4*(1-0.38)
= 2.48
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=2.48*0.5+11.03*0.5
WACC =6.76%

d

D/A = 0.3

D/E = D/(A-D) = 0.3/(1-0.3)=0.4285

Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 8.34+0.4285*(8.34-4)*(1-0.38)
Levered cost of equity = 9.49
Weight of equity = 1-D/A
Weight of equity = 1-0.3
W(E)=0.7
Weight of debt = D/A
Weight of debt = 0.3
W(D)=0.3
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 4*(1-0.38)
= 2.48
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=2.48*0.3+9.49*0.7
WACC =7.39%